OpenStudy (anonymous):

Find the mean of the following group of numbers: 6.1, 5.9, 4.2, 5.9, 10.7. Round your answer to one decimal place. Answer

OpenStudy (anonymous):

just add all together and divide it by 5

OpenStudy (anonymous):

\[\frac{6.1 + 5.9 + 4.2 +5.9 + 10.7}{5} = \frac{32.8}{5} = 6.56 \approx 6.6\]